OpenStudy (anonymous):

• What are the differences among expressions, equations, and functions? Provide examples of each.

OpenStudy (anonymous):

you here??

OpenStudy (anonymous):

okay functions is the simple terms is where put in one thing and it comes out anther. F(X) if x=1 what is the function of x. an expression is something like 2x+4=5, equations want you to solve for a certain letter.

OpenStudy (anonymous):

hello??

OpenStudy (rsadhvika):

expression : 2x, 3x-1 etc equation : y^2 = 4x function : y = x^2 only difference between an equation and a function is that, an equation need not pass the vertical line test. all functions are equations. but not all equations are functions. only some equations are functions, and, all equations have expression on each side. HTH

OpenStudy (anonymous):

okay I admit made it harder than it had to be.